# A family of log-correlated Gaussian processes

### v1

@@ 4.1. A stochastic-integral representation
$\mathbb{P}^{\prime}_{r}$ $\left(m(A_{x}){\rm~{}is~{}odd},m(A_{y}){\rm~{}is~{}odd}\right)$
$=\mathbb{P}^{\prime}_{r}\left(m(A_{x}\setminus A_{y}){\rm~{}is~{}odd}\right)\mathbb{P}^{\prime}_{r}\left(m(A_{y}\setminus A_{x}){\rm~{}is~{}odd}\right)\mathbb{P}^{\prime}_{r}\left(m(A_{x}\cap A_{y}){\rm~{}is~{}even}\right)$
$\quad+\mathbb{P}^{\prime}_{r}\left(m(A_{x}\setminus A_{y}){\rm~{}is~{}even}\right)\mathbb{P}^{\prime}_{r}\left(m(A_{y}\setminus A_{x}){\rm~{}is~{}even}\right)\mathbb{P}^{\prime}_{r}\left(m(A_{x}\cap A_{y}){\rm~{}is~{}odd}\right)$
$=\frac{1}{4}\left(1-e^{-2r\mu(A_{x})}-e^{-2r\mu(A_{y})}+e^{-2r\mu(A_{x}\Delta A_{y})}\right).$
Then, taking $\widetilde{\Gamma}_{r}(x,y)$ as given in (4.4), we have
(4.7) $\widetilde{\Gamma}_{r}(x,y)$ $=\mathbb{P}^{\prime}_{r}\left(m(A_{x}){\rm~{}is~{}odd},m(A_{y}){\rm~{}is~{}odd}\right)-\mathbb{P}^{\prime}_{r}\left(m(A_{x}){\rm~{}is~{}odd}\right)\mathbb{P}^{\prime}_{r}\left(m(A_{y}){\rm~{}is~{}odd}\right)$
$=\frac{1}{4}\left(e^{-2r\mu(A_{x}\Delta A_{y})}-e^{-2r(\mu(A_{x})+\mu(A_{y}))}\right)\geq 0.$
@@
$\Psi_{2,f}(r)$ $={\mathbb{E}}^{\prime}_{r}\int_{\mathsf{M}}\int_{\mathsf{M}}\left({{\bf 1}}_{\left\{m(A_{x}){\rm~{}is~{}odd}\right\}}-\widetilde{p}_{r}(A_{x})\right)\left({{\bf 1}}_{\left\{m(A_{y}){\rm~{}is~{}odd}\right\}}-\widetilde{p}_{r}(A_{y})\right)f(x)f(y)\lambda({\rm d}x)\lambda({\rm d}y)$
$=\int_{\mathsf{M}}\int_{\mathsf{M}}\widetilde{\Gamma}_{r}(x,y)f(x)f(y)\lambda({\rm d}x)\lambda({\rm d}y)$
$=\frac{1}{4}\int_{\mathsf{M}}\int_{\mathsf{M}}\left(e^{-2r\mu(A_{x}\Delta A_{y})}-e^{-2r(\mu(A_{x})+\mu(A_{y})}\right)f(x)f(y)\lambda({\rm d}x)\lambda({\rm d}y).$
@@
For the sake of simplicity, assume that $f\geq 0$. Write
$e^{-2r\mu(A_{x}\Delta A_{y})}$ $-e^{-2r(\mu(A_{x})+\mu(A_{y}))}$
(4.8) $=e^{-2r\mu(A_{x}\Delta A_{y})}\left(1-e^{-4r\mu(A_{x}\cap A_{y})}\right)$
(4.9) $=e^{-2r\mu(A_{x})}\left(e^{-2r(\mu(A_{y}\setminus A_{x})-\mu(A_{x}\cap A_{y}))}-e^{-2r\left(\mu(A_{y}\setminus A_{x})+\mu(A_{x}\cap A_{y})\right)}\right).$
Using (4.8), we write
$\Psi_{2,f}(r)$ $\leq\frac{1}{2}\int_{\mathsf{M}}\int_{{y\in{\mathsf{M}}:\mu(A_{y})\geq\mu(A_{x})}}f(x)f(y)e^{-2r\mu(A_{x}\Delta A_{y})}\left(1-e^{-4r\mu(A_{x}\cap A_{y})}\right)\lambda({\rm d}x)\lambda({\rm d}y),$
and break the outer integral into two parts. We have
$\int_{\mu(A_{x})>1/r}$ $\int_{\mu(A_{y})\geq\mu(A_{x})}f(x)f(y)e^{-2r\mu(A_{x}\Delta A_{y})}\left(1-e^{-4r\mu(A_{x}\cap A_{y})}\right)\lambda({\rm d}x)\lambda({\rm d}y)$
$\leq\int_{\mu(A_{x})>1/r}f(x)\int_{\mu(A_{y})\geq\mu(A_{x})}f(y)\lambda({\rm d}y)\lambda({\rm d}x)\leq\int_{\mu(A_{x})>1/r}f(x)\lambda({\rm d}x){\left\|f\right\|}_{1}$
$\leq r^{\delta}{\left\|f\right\|}_{1}\int_{\mu(A_{x})>1/r}\mu(A_{x})^{\delta}f(x)\lambda({\rm d}x)\leq C{\left\|f\right\|}_{1}{\left\|f\right\|}_{1,\delta}r^{\delta},$
and
$\int_{\mu(A_{x})\leq 1/r}\int_{\mu(A_{y})\geq\mu(A_{x})}f(x)f(y)e^{-2r\mu(A_{x}\Delta A_{y})}\left(1-e^{-4r\mu(A_{x}\cap A_{y})}\right)\lambda({\rm d}x)\lambda({\rm d}y)$
$\quad\leq\int_{\mu(A_{x})\leq 1/r}4r\mu(A_{x}\cap A_{y})f(x)\int_{\mu(A_{y})\geq\mu(A_{x})}f(y)\lambda({\rm d}y)\lambda({\rm d}x)$
$\quad\leq{\left\|f\right\|}_{1}\int_{\mu(A_{x})\leq 1/r}4r\mu(A_{x})f(x)\lambda({\rm d}x)\leq 4{\left\|f\right\|}_{1}r^{\delta}\int_{\mu(A_{x})\leq 1/r}\mu(A_{x})^{\delta}f(x)\lambda({\rm d}x)\leq Cr^{\delta}.$
Next, using (4.9) and bounding its right-hand side by $e^{-2r\mu(A_{x})}$, we have
$|\Psi_{2,f}(r)|\leq\int_{\mathsf{M}}e^{-2r\mu(A_{x})}f(x)\int_{\mathsf{M}}f(y)\lambda({\rm d}y)\lambda({\rm d}x)\leq{\left\|f\right\|}_{1}\int_{\mathsf{M}}e^{-2r\mu(A_{x})}f(x)\lambda({\rm d}x)\leq Cr^{-\delta^{\prime}},$
where the last step follows from (iii) of Assumption 4.1. We have thus proved
(4.10) $\Psi_{2,f}(r)\leq C(r^{\delta}\wedge r^{-\delta^{\prime}})\mbox{ for all }r>0,$
for all $f\in{\mathcal{F}}_{0}({\mathsf{M}})$ (the constant $C$ depends on $f$). Now (4.6) follows with $\alpha=2$. For $\alpha\in(0,2)$, using $\Psi_{\alpha,f}(r)\leq\Psi_{2,f}(r)^{\alpha/2}$, (4.6) remains to hold. We have thus proved that (4.1) holds and the right-hand side is finite.
In particular for the case $\alpha=2$, (4.1) yields
${\rm{Cov}}(G_{2}(f),G_{2}(f))$ $=4\int_{0}^{\infty}\int_{\mathsf{M}}\int_{\mathsf{M}}f(x)f(y)\widetilde{\Gamma}_{r}(x,y)\lambda({\rm d}x)\lambda({\rm d}y)\frac{{\rm d}r}{r}$
$=\int_{\mathsf{M}}\int_{\mathsf{M}}f(x)f(y)\int_{0}^{\infty}\frac{e^{-2r\mu(A_{x}\Delta A_{y})}-e^{-2r(\mu(A_{x})+\mu(A_{y}))}}{r}{\rm d}r\lambda({\rm d}x)\lambda({\rm d}y)$
$=\int_{\mathsf{M}}\int_{\mathsf{M}}f(x)f(y)\log\frac{\mu(A_{x})+\mu(A_{y})}{\mu(A_{x}\Delta A_{y})}\lambda({\rm d}x)\lambda({\rm d}y),$
@@
This time, we can relate $G_{\alpha}$ to an extension of the decomposition of Lei and Nualart (see [14, Section 3.2]). Given a metric space $({\mathsf{M}},{\mathsf{d}})$ and $H\in(0,1/2)$, we have
(4.11) $\left(X_{\alpha}^{H}(x)\right)_{x\in{\mathsf{M}}}\stackrel{{\scriptstyle d}}{{=}}\left(G_{\alpha}^{H}(x)+W_{\alpha}^{H}(x)\right)_{x\in{\mathsf{M}}}$
with
$X_{\alpha}^{H}(x)$ :=+×$:=\int_{\mathbb{R}_{+}\times{\mathbb{M}}_{p}(E)}{{\bf 1}}_{\left\{m(A_{x}){\rm~{}is~{}odd}\right\}}M_{\alpha,H}({\rm d}r{\rm d}m),$
$G_{\alpha}^{H}(x)$ :=+×$:=\int_{\mathbb{R}_{+}\times{\mathbb{M}}_{p}(E)}\left({{\bf 1}}_{\left\{m(A_{x}){\rm~{}is~{}odd}\right\}}-\widetilde{p}_{r}(x)\right)M_{\alpha,H}({\rm d}r{\rm d}m),$
$W_{\alpha}^{H}(x)$ :=+×$:=\int_{\mathbb{R}_{+}\times{\mathbb{M}}_{p}(E)}\widetilde{p}_{r}(x)M_{\alpha,H}({\rm d}r{\rm d}m),$
where $M_{\alpha,H}$ is a symmetric $\alpha$-stable random measure on +×$\mathbb{R}_{+}\times\mathbb{M}_{p}(E)$ with control measure $C_{2H}r^{-2H-1}{\rm d}r{\rm d}\mathbb{P}^{\prime}_{r}$ with $C_{\beta}=\beta 2^{-\beta}/\Gamma(1-\beta),\beta\in(0,1)$. When $\alpha=2$, (4.11) is the decomposition of fractional Brownian motion with Hurst index $H$ on $({\mathsf{M}},{\mathsf{d}})$. The covariance function is related to the measure definite kernel ${\mathsf{d}}$ as
(4.12) ${\rm{Cov}}\left(G_{2}^{H}(x),G_{2}^{H}(y)\right)=\left(\mu(A_{x})+\mu(A_{y})\right)^{2H}-\mu(A_{x}\Delta A_{y})^{2H}.$
With $\alpha\in(0,2)$, (4.11) can be viewed as an extension to stable processes (although this time $G_{\alpha}^{H}$ and $W_{\alpha}^{H}$ are no longer independent). More details can be found in [14], and these processes are closely related to set-indexed fractional Brownian motions [19]. We have the following.
Lemma 4.3.
With $G_{\alpha}^{H}(f):=\int_{\mathsf{M}}G_{\alpha}^{H}(x)f(x)\lambda({\rm d}x)$, we have
$\frac{1}{H^{1/\alpha}}G_{\alpha}^{H}(f)\Rightarrow G_{\alpha}(f)$
as $H\downarrow 0$ for all $f\in{\mathcal{F}}_{0}({\mathsf{M}})$.
The proof is by computing the characteristic functions. The convergence of covariance kernel follows immediately from (4.12). All these are straightforward and omitted.
@@ 4.2. A central limit theorem
We consider another aggregated model combining ideas from [11] and [14].
@@
for all $f\in{\mathcal{F}}_{0}({\mathsf{M}})$. This represents an aggregated model based on the triangular arrays of i.i.d. point processes $(N_{n,i}^{(q_{i})})_{n\in\mathbb{N},i\in\mathbb{N}}$.
@@
$\lim_{n\to\infty}{\mathbb{E}}\exp\left({\rm i}G_{\alpha,n}(f)\right)={\mathbb{E}}\exp\left({\rm i}G_{\alpha}(f)\right),\mbox{ for all }f\in{\mathcal{F}}_{0}({\mathsf{M}}).$
@@
and (recall $\Psi_{\alpha,f}(r)$ in (4.5))
(4.16) ${\mathbb{E}}\widehat{\Psi}_{\alpha,f}(r)=\Psi_{\alpha,f}(r)\leq\left(\Psi_{2,f}(r)\right)^{\alpha/2}\leq C(r^{\delta}\wedge r^{-\delta^{\prime}})^{\alpha/2},$
for some constant $C$ depending on $f\in{\mathcal{F}}_{0}({\mathsf{M}})$, where the last step was shown in (4.10). Now, combining (4.14), (4.15) and (4.16) and letting $\epsilon\downarrow 0$, the above estimates yield (4.13). ∎
@@ Appendix A An example by Iksanov and Kotelnikova
Iksanov and Kotelnikova, [22] considered a limit theorem for the Karlin model in the $\alpha=0$ regime. We omit the description of the model (they actually investigated a sophisticated hierarchical Karlin model and the discussions here concern only their results at the first level), but focus on the limit processes. In particular, they investigated the scaling limit of the $j$-occupancy process. That is, the counting process for the number of urns containing exactly $j$ after $n$ rounds. The limit process [22, Eq. (4)] is denoted by $(X_{j}(u))_{u\in\mathbb{R}}$ below. The parameterization is a little unconventional as the index set is $\mathbb{R}$, although we shall eventually see that applying a time-change $u\mapsto t=\log u\in(0,\infty)$ we can relate $(X_{j})_{j\in\mathbb{N}}$ to our process.
Stochastic-integral representations for their limit processes were provided in [22]. Some preparations are needed for the Gaussian random measures involved. Let $\mathbb{P}_{U^{\infty}}$ denote the uniform probability measure on $[0,1]^{\mathbb{N}}$ (the distribution of i.i.d. uniform random variables). Let $W_{\infty}$ denote a Gaussian random measure on ×$\mathbb{R}\times[0,1]^{\mathbb{N}}$ with intensity measure ${\rm d}x{\rm d}\mathbb{P}_{U^{\infty}}$. For each $j\in\mathbb{N}$, let $W_{j}$ denote $W_{\infty}$ restricted to ×$\mathbb{R}\times[0,1]^{j}$, that is,
@@
(A.2) $\left(X_{j}(\log t)\right)_{t\in(0,\infty),j\in\mathbb{N}}\stackrel{{\scriptstyle d}}{{=}}\left(\int_{\mathbb{R}\times\Omega^{\prime}}\left({{\bf 1}}_{\left\{N^{\prime}(rt)=j\right\}}-\mathbb{P}(N^{\prime}(rt)=j)\right)M({\rm d}r{\rm d}\omega^{\prime})\right)_{t\in(0,\infty),j\in\mathbb{N}}.$
@@
$\left\{\prod_{k=1}^{j+1}Y_{k}<\exp\left(-e^{-(x-u)}\right)<\prod_{k=1}^{j}Y_{k}\right\}$ =$=\left\{\sum_{i=1}^{j+1}(-\log Y_{i})>e^{-(x-u)}>\sum_{i=1}^{j}(-\log Y_{i})\right\}$
=$=\left\{\sum_{i=1}^{j}E_{i}<e^{-(x-u)}<\sum_{i=1}^{j+1}E_{i}\right\}=\left\{N(e^{-(x-u)})=j\right\},$
@@
Then, with $\widetilde{X}_{j}(t)=X_{j}(\log t),t\in(0,\infty)$, $\sum_{\ell=1}^{\infty}\widetilde{X}_{2j-1}(t)$ is formally (since this is not point-wise convergent) our generalized Gaussian process: compare with $G(f)$ in (2.2) (up to a multiplicative constant as the control measures differ by a multiplicative constant $2$).

### v3

@@ 4.1. A stochastic-integral representation
$\mathbb{P}^{\prime}_{\beta,r}$ $\left(m(A_{x}){\rm~is~odd},m(A_{y}){\rm~is~odd}\right)$
$=\mathbb{P}^{\prime}_{\beta,r}\left(m(A_{x}\setminus A_{y}){\rm~is~odd}\right)\mathbb{P}^{\prime}_{\beta,r}\left(m(A_{y}\setminus A_{x}){\rm~is~odd}\right)\mathbb{P}^{\prime}_{\beta,r}\left(m(A_{x}\cap A_{y}){\rm~is~even}\right)$
$\quad+\mathbb{P}^{\prime}_{\beta,r}\left(m(A_{x}\setminus A_{y}){\rm~is~even}\right)\mathbb{P}^{\prime}_{\beta,r}\left(m(A_{y}\setminus A_{x}){\rm~is~even}\right)\mathbb{P}^{\prime}_{\beta,r}\left(m(A_{x}\cap A_{y}){\rm~is~odd}\right)$
$={\mathbb{E}}^{\prime}_{\beta,r}\left(\frac{1}{4}\left(1-e^{-2\mathcal{S}_{\beta,r}\mu(A_{x})}-e^{-2\mathcal{S}_{\beta,r}\mu(A_{y})}+e^{-2\mathcal{S}_{\beta,r}\mu(A_{x}\Delta A_{y})}\right)\right)$
$=\frac{1}{4}\left(1-e^{-(2\mu(A_{x}))^{\beta}r}-e^{-(2\mu(A_{y}))^{\beta}r}+e^{-(2\mu(A_{x}\Delta A_{y}))^{\beta}r}\right).$
Then, taking $\Gamma^{(\beta)}_{r}(x,y)$ as given in (4.4), we have
(4.7) $\Gamma_{r}^{(\beta)}(x,y)$ $=\mathbb{P}^{\prime}_{\beta,r}\left(m(A_{x}){\rm~is~odd},m(A_{y}){\rm~is~odd}\right)-\mathbb{P}^{\prime}_{\beta,r}\left(m(A_{x}){\rm~is~odd}\right)\mathbb{P}^{\prime}_{\beta,r}\left(m(A_{y}){\rm~is~odd}\right)$
$=\frac{1}{4}\left(e^{-(2\mu(A_{x}\Delta A_{y}))^{\beta}r}-e^{-((2\mu(A_{x}))^{\beta}+(2\mu(A_{y}))^{\beta})r}\right)\geq 0.$
@@
$\Psi_{2,\beta,f}(r)$
$={\mathbb{E}}^{\prime}_{\beta,r}\int_{\mathsf{M}}\int_{\mathsf{M}}\left({{\bf 1}}_{\left\{m(A_{x}){\rm~is~odd}\right\}}-\widetilde{p}_{\beta,r}(A_{x})\right)\left({{\bf 1}}_{\left\{m(A_{y}){\rm~is~odd}\right\}}-\widetilde{p}_{\beta,r}(A_{y})\right)f(x)f(y)\lambda({\rm d}x)\lambda({\rm d}y)$
$=\int_{\mathsf{M}}\int_{\mathsf{M}}\Gamma_{r}^{(\beta)}(x,y)f(x)f(y)\lambda({\rm d}x)\lambda({\rm d}y)$
$=\frac{1}{4}\int_{\mathsf{M}}\int_{\mathsf{M}}\left(e^{-(2\mu(A_{x}\Delta A_{y}))^{\beta}r}-e^{-((2\mu(A_{x}))^{\beta}+(2\mu(A_{y}))^{\beta})r}\right)f(x)f(y)\lambda({\rm d}x)\lambda({\rm d}y).$
@@
Next, we prove
(4.8) $\Psi_{2,\beta,f}(r)\leq C(r^{\delta}\wedge r^{-\delta^{\prime}})\mbox{ for all }r>0,$
for all $f\in{\mathcal{F}}_{\delta,\delta^{\prime},\beta}({\mathsf{M}})$ (the constant $C$ depends on $f$). For the sake of simplicity, assume that $f\geq 0$. We break the outer integral into two parts and proceed as follows. First,
$\int_{\mu(A_{x})>1/r}$ $\int_{\mathsf{M}}f(x)f(y)\left(e^{-(2\mu(A_{x}\Delta A_{y}))^{\beta}r}-e^{-((2\mu(A_{x}))^{\beta}+(2\mu(A_{y}))^{\beta})r}\right)\lambda({\rm d}x)\lambda({\rm d}y)$
$\leq\int_{\mu(A_{x})>1/r}\int_{\mathsf{M}}f(x)f(y)\lambda({\rm d}x)\lambda({\rm d}y)=\int_{\mu(A_{x})>1/r}f(x)\lambda({\rm d}x){\left\|f\right\|}_{1}$
$\leq r^{\delta}{\left\|f\right\|}_{1}\int_{\mu(A_{x})>1/r}\mu(A_{x})^{\delta}f(x)\lambda({\rm d}x)\leq C{\left\|f\right\|}_{1}{\left\|f\right\|}_{1,\delta}r^{\delta}.$
Second, using
$e^{-(2\mu(A_{x}\Delta A_{y}))^{\beta}r}-e^{-((2\mu(A_{x}))^{\beta}+(2\mu(A_{y}))^{\beta})r}$ $\leq 1-e^{-2^{\beta}(\mu^{\beta}(A_{x})+\mu^{\beta}(A_{y})-\mu^{\beta}(A_{x}\Delta A_{y}))r}$
$\leq 2^{\beta}(\mu^{\beta}(A_{x})+\mu^{\beta}(A_{y})-\mu^{\beta}(A_{x}\Delta A_{y}))r$
$\leq 2^{\beta+1}\mu^{\beta}(A_{x})r,$
we have
$\int_{\mu(A_{x})\leq 1/r}$ $\int_{\mathsf{M}}f(x)f(y)\left(e^{-(2\mu(A_{x}\Delta A_{y}))^{\beta}r}-e^{-((2\mu(A_{x}))^{\beta}+(2\mu(A_{y}))^{\beta})r}\right)\lambda({\rm d}x)\lambda({\rm d}y)$
$\leq{\left\|f\right\|}_{1}\int_{\mu(A_{x})\leq 1/r}2^{\beta+1}r\mu^{\beta}(A_{x})f(x)\lambda({\rm d}x)$
$\leq Cr\int_{\mu(A_{x})\leq 1/r}(1/r)^{\beta(1-\widetilde{\delta})}\mu^{\beta\widetilde{\delta}}(A_{x})f(x)\lambda({\rm d}x)\leq Cr^{1-\beta+\beta\widetilde{\delta}}{\left\|f\right\|}_{1,\beta\widetilde{\delta}}$,
for $\widetilde{\delta}\in(0,1)$. Taking $\widetilde{\delta}=\delta/\beta$, the above is bounded by $Cr^{1-\beta+\delta}\leq Cr^{\delta}$ for $r\in[0,1]$. We have thus proved $\Psi_{2,\beta,f}(r)\leq Cr^{\delta}$ for $r\in[0,1]$. Alternatively, using $e^{-(2\mu(A_{x}\Delta A_{y}))^{\beta}r}-e^{-((2\mu(A_{x}))^{\beta}+(2\mu(A_{y}))^{\beta})r}\leq e^{-2r\mu(A_{x}\Delta A_{y})}=e^{-2^{\beta}{\mathsf{d}}^{\beta}(x,y)r}$, we have
$|\Psi_{2,\beta,f}(r)|\leq\int_{\mathsf{M}}\int_{\mathsf{M}}e^{-2^{\beta}{\mathsf{d}}^{\beta}(x,y)r}f(x)f(y)\lambda({\rm d}y)\lambda({\rm d}x)\leq Cr^{-\delta^{\prime}},$
where the last step follows from (iii) of Assumption 4.1. We have thus proved (4.8).
Now (4.6) follows with $\alpha=2$. For $\alpha\in(0,2)$, using $\Psi_{\alpha,\beta,f}(r)\leq\Psi_{2,\beta,f}(r)^{\alpha/2}$, (4.6) remains to hold. We have thus proved that (4.1) holds and the right-hand side is finite.
In particular for the case $\alpha=2$, (4.1), (4.7) and Gamma integral yield
${\rm{Cov}}(G_{2,\beta}(f),G_{2,\beta}(f))$ $=4\int_{0}^{\infty}\int_{\mathsf{M}}\int_{\mathsf{M}}f(x)f(y)\Gamma_{r}^{(\beta)}(x,y)\lambda({\rm d}x)\lambda({\rm d}y)\frac{{\rm d}r}{r}$
$=\int_{\mathsf{M}}\int_{\mathsf{M}}f(x)f(y)\log\frac{\mu^{\beta}(A_{x})+\mu^{\beta}(A_{y})}{\mu^{\beta}(A_{x}\Delta A_{y})}\lambda({\rm d}x)\lambda({\rm d}y),$
@@ 4.2. A central limit theorem
We consider another aggregated model combining ideas from [11] and [14] that scales to the generalized process just introduced. We only illustrate the case $H=1/2$.
@@
for all $f\in{\mathcal{F}}_{\delta,\delta^{\prime},1}({\mathsf{M}})$. This represents an aggregated model based on the triangular arrays of i.i.d. point processes $(N_{n,i}^{(q_{i})})_{n\in\mathbb{N},i\in\mathbb{N}}$.
@@
$\lim_{n\to\infty}{\mathbb{E}}\exp\left({\rm i}G_{\alpha,n}(f)\right)={\mathbb{E}}\exp\left({\rm i}G_{\alpha}(f)\right),\mbox{ for all }f\in{\mathcal{F}}_{\delta,\delta^{\prime},1}({\mathsf{M}}).$
@@
and (recall $\Psi_{\alpha,1,f}(r)$ in (4.5))
(4.12) ${\mathbb{E}}\widehat{\Psi}_{\alpha,f}(r)=\Psi_{\alpha,1,f}(r)\leq\left(\Psi_{2,1,f}(r)\right)^{\alpha/2}\leq C(r^{\delta}\wedge r^{-\delta^{\prime}})^{\alpha/2},$
for some constant $C$ depending on $f\in{\mathcal{F}}_{\delta,\delta^{\prime},1}({\mathsf{M}})$, where the last step was shown in (4.8). Now, combining (4.10), (4.11) and (4.12) and letting $\epsilon\downarrow 0$, the above estimates yield (4.9). ∎
@@ Appendix A An example by Iksanov and Kotelnikova
Iksanov and Kotelnikova, [22] considered a limit theorem for the Karlin model in the regime that the sampling frequencies $(p_{j})_{j\in\mathbb{N}}$ decay faster than any polynomial rate (recall that for the Karlin model with $\alpha\in(0,1)$, $p_{j}\sim Cj^{-1/\alpha}$). Namely, they considered the case that for $\rho(t):=|\{k\in\mathbb{N}:p_{k}\geq 1/t\}|$, that there exists a slowly varying function $g$ at infinity with $g(x)\to\infty$ as $x\to\infty$ such that $\lim_{x\to\infty}(\rho(\lambda t)-\rho(t))/g(t)=\log\lambda$ for all $\lambda>0$. We omit the description of the model (they actually investigated a sophisticated hierarchical Karlin model and the discussions here concern only their results at the first level), but focus on the limit processes.
In particular, they investigated the scaling limit of the $j$-occupancy process. That is, the counting process for the number of urns containing exactly $j$ after $n$ rounds. The limit process [22, Eq. (4)] is denoted by $(X_{j}(u))_{u\in\mathbb{R}}$ below. The parameterization is a little unconventional as the index set is $\mathbb{R}$, although we shall eventually see that applying a time-change $u\mapsto t=\log u\in(0,\infty)$ we can relate $(X_{j})_{j\in\mathbb{N}}$ to the process of interest here $G^{1/2}$.
Stochastic-integral representations for their limit processes were provided in [22]. Some preparations are needed for the Gaussian random measures involved. Let $\mathbb{P}_{U^{\infty}}$ denote the uniform probability measure on $0,1]^{\mathbb{N}}$ (the distribution of i.i.d. uniform random variables). Let $W_{\infty}$ denote a Gaussian random measure on $\mathbb{R}\times[0,1]^{\mathbb{N}}$ with intensity measure ${\rm d}x{\rm d}\mathbb{P}_{U^{\infty}}$. For each $j\in\mathbb{N}$, let $W_{j}$ denote $W_{\infty}$ restricted to $\mathbb{R}\times[0,1]^{j}$, that is,
@@
$\left(X_{j}(\log t)\right)_{t\in(0,\infty),j\in\mathbb{N}}\stackrel{{\scriptstyle d}}{{=}}\left(\int_{\mathbb{R}\times\Omega^{\prime}}\left({{\bf 1}}_{\left\{N^{\prime}(rt)=j\right\}}-\mathbb{P}(N^{\prime}(rt)=j)\right)M({\rm d}r{\rm d}\omega^{\prime})\right)_{t\in(0,\infty),j\in\mathbb{N}}.$
@@
$\left\{\prod_{k=1}^{j+1}Y_{k}<\exp\left(-e^{-(x-u)}\right)<\prod_{k=1}^{j}Y_{k}\right\}$ $=\left\{\sum_{i=1}^{j+1}(-\log Y_{i})>e^{-(x-u)}>\sum_{i=1}^{j}(-\log Y_{i})\right\}$
$=\left\{\sum_{i=1}^{j}E_{i}<e^{-(x-u)}<\sum_{i=1}^{j+1}E_{i}\right\}=\left\{N(e^{-(x-u)})=j\right\},$
@@
Then, with $\widetilde{X}_{j}(t)=X_{j}(\log t),t\in(0,\infty)$, $\sum_{\ell=1}^{\infty}\widetilde{X}_{2j-1}(t)$ is formally (since this is not point-wise convergent) our generalized Gaussian process: compare with $G^{1/2}(f)$ in (2.2) (up to a multiplicative constant as the control measures differ by a multiplicative constant $2$).
Processes with similar representations often arise as scaling limits of counting processes associated to models with certain renewal structures. See Alsmeyer et al., [1] for a limit theorem on the counting process for a family of decoupled renewal processes.